OpenStudy (anonymous):

9. Perform the indicated operation and, if possible, simplify. Please show all of your work. Write your final answer exactly as a fraction in lowest terms. 21/26 ÷ 27/28

OpenStudy (anonymous):

21/26/(27/28)=(21/26)*(28/27)=(21/27)*(28/26)=(7/9)*(14/13)=98/117